# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ A.2. Scenario: Student
In this scenario, the story prefix includes a team member with a traditionally feminine (“Abby”) or masculine (“John”) name. The model then suggests that this character is “competitive” or “unassertive” (see examples in Table 5). According to Cao et al. (2022b), American annotators view men as comparatively more competitive and women as comparatively more unassertive.
@@ A.2.1. Effects on Competitiveness Disaggregated by Gender
When we consider competitiveness rates in overall stories, we see that suggestions often increase the respective rates. We see significantly more unassertive masc-coded ($t(117)=-3.024$, $p_{\text{FDR}}\approx 0.0157$), unassertive fem-coded ($t(116)=-2.874$, $p_{\text{FDR}}\approx 0.0210$), and competitive fem-coded students ($t(125)=-3.464$, $p_{\text{FDR}}\approx 0.0044$) when suggested when comparing to no suggestions conditions. However, due to a higher baseline rate of competitive masc-coded students in the no suggestions conditions, we do not see a significant effect when providing competitive masc-coded suggestions ($t(121)=-1.297$, $p_{\text{FDR}}\approx 0.3747$). This provides some evidence toward H1 as the pro-stereotypical competitive masc-coded condition is similar to behavior without suggestions.
We also see that generally providing suggestions increases competitiveness vs unassertiveness rates within a gender. For example, providing competitive masc-coded suggestions significantly increases the rate of competitive masc-coded students over unassertive masc-coded students ($t(174)=-3.898$, $p_{\text{FDR}}\approx 0.0010$). We see similar significant trends for unassertive masc-coded students ($t(166)=2.872$, $p_{\text{FDR}}\approx 0.0207$) and competitive fem-coded students ($t(174)=-8.131$, $p_{\text{FDR}}<0.0001$). However, we see no significant difference with unassertive fem-coded suggestions ($t(156)=0.883$, $p_{\text{FDR}}\approx 0.5725$). This provides some evidence against H1, implying that participants may be more resistant to unassertive fem-coded student suggestions despite existing stereotypes of gender and competitiveness.
@@
When we consider the overall rate of reliance on model suggestions, we see that participants typed new words or edited model suggestions significantly more in the unassertive fem-coded student condition than the unassertive masc-coded student condition ($t(3615)=4.453$, $p_{\text{FDR}}\approx 0.0001$) and similarity for the competitive masc-coded student condition over the unassertive masc-coded student condition ($t(4135)=-6.49$, $p_{\text{FDR}}<0.0001$). These results do not support H2, but we see that none of these trends are significantly present when we constrain to words that determine the character’s competitiveness (H2a). When we consider model suggestions that affect competitiveness, we see marginally more acceptance of competitive fem-coded suggestions than unassertive (H2b; $t(426)=-2.159$, $p_{\text{FDR}}\approx 0.0999$).
Overall, the results in this scenario were mixed. We find a relatively low overall rate of clearly specifying the given student as competitive or unassertive which may have skewed the results. This may be due to poor scenario design where participants decided to focus on topics other than leadership or competitiveness, or it could be that in the classroom settings, participants held weaker internal stereotypes about gender and competitiveness than Cao et al. (2022b)’s held about more general settings.
@@ A.3. Scenario: Teachers
In this scenario, participants write about a teacher with a given gender where the model attempts to suggest the teacher’s personality as “likeable” or “repellent” (see examples in Table 5). According to Cao et al. (2022b), American annotators view men as comparatively less likeable than women.
A.3.1. Effects on Likeability Disaggregated by Gender
Here, we discuss how suggestions affect the likeability stance of “Mrs. Brown” and “Mr. Brown” (Table 4).
@@ A.3.1. Effects on Likeability Disaggregated by Gender
Considering overall stories, we see that regardless of the teacher’s presumed gender, participants made the teacher likeable significantly more often than repellent when not given suggestions (fem-coded teacher: $t(68)=-4.747$, $p_{\text{FDR}}\approx 0.0001$; masc-coded teacher: $t(74)=-3.195$, $p_{\text{FDR}}\approx 0.0110$). These trends continue to be significant regardless of suggestion type, including repellent suggestions (likeable fem-coded: $t(162)=-12.726$, $p_{\text{FDR}}<0.0001$; repellent fem-coded: $t(150)=-3.886$, $p_{\text{FDR}}\approx 0.0011$; likeable masc-coded: $t(184)=-8.86$, $p_{\text{FDR}}<0.0001$; repellent masc-coded: $t(166)=-3.774$, $p_{\text{FDR}}\approx 0.0016$). In other words, participants preferred to make the teacher likeable, regardless of the presence or type of suggestions and regardless of the teacher’s gender.
We also find that “Mr. Brown” is written as likable marginally more often with likeable suggestions than without suggestions ($t(129)=-2.489$, $p_{\text{FDR}}\approx 0.0516$). However, perhaps due to a higher base rate of “Mrs. Brown” being written as likeable, we see no such increase in likeability with added likeable suggestions for “Mrs. Brown” ($t(115)=-1.403$, $p_{\text{FDR}}\approx 0.3291$). In other words participants may have a stronger default preference for “Mrs. Brown” being likeable, leading to a more limited effect of likeable suggestions. This is in line with Cao et al. (2022b)’s findings about gender-likeability stereotypes in humans. And these findings provide some support for H1 in that the proportion of likeable “Mrs. Brown”s (pro-stereotypical) with no suggestions is more similar to the proportion of likeable “Mrs. Brown”s with likeable suggestions than the the proportion of likeable “Mr. Brown”s (anti-stereotypical) with no suggestions is to the proportion of likeable “Mr. Brown”s with likeable suggestions.
@@
When we consider the rate of acceptances of model suggestions, we see significantly less acceptance of model suggestions in the condition where “Mrs. Brown” is suggested to be repellent over likeable ($t(3591)=2.618$, $p_{\text{FDR}}\approx 0.0351$), while we see a significant effect in the opposite direction for “Mr. Brown” ($t(3892)=-3.162$, $p_{\text{FDR}}\approx 0.0086$). We also see significantly less acceptance in the “Mr. Brown” is likeable condition than for “Mrs. Brown” ($t(4013)=-5.922$, $p_{\text{FDR}}<0.0001$). These results support H2, as we can see that more pro-stereotypical conditions (e.g., suggesting a fem-coded character is likeable) lead to more acceptance of suggestions.
However, the effects are quite different when we only consider words that determine likeability. Here we see a trend of participants accepting more “likeable” suggestions over “repellent” for either gender, though the effect is only significant for “Mr. Brown” (masc-coded: $t(179)=2.673$, $p_{\text{FDR}}\approx 0.0331$; fem-coded: $t(157)=2.068$, $p_{\text{FDR}}\approx 0.1190$). Under H2a, we would have expected to see less acceptance of “likeable” suggestions for the masc-coded “Mr. Brown”. When we consider the acceptance rate of model suggestions, we see higher rates of acceptance of “likeable” suggestions for either gender, but in this case, it is only significant for “Mrs. Brown” (masc-coded: $t(438)=-2.446$, $p_{\text{FDR}}\approx 0.0533$, $p_{\text{FDR}}\approx 0.0331$; fem-coded: $t(383)=-4.106$, $p_{\text{FDR}}\approx 0.0004$), supporting H2b.
These results show that participants may have preferred suggestions of teachers of any gender being likeable. This seems reasonable as at the story-level likeable teachers were generally preferred even without suggestions. While these results do not match our overall hypotheses about reliance under pairs of gender and likeability, they may suggest that participants’ stereotypes about teachers being likeable people were stronger than their stereotypes about people of different genders being likeable.
@@ A.4. Scenario: Town Hall
In this scenario, the story prefix includes a town hall participant with a traditionally feminine (“Rebecca”) or masculine (“Thomas”) name. The town hall is about an affordable housing development, and the model suggests that the character has a conservative or liberal viewpoint on this issue (see examples in Table 5). According to Cao et al. (2022b), American annotators view men as comparatively more conservative than women.
@@ A.4.1. Effects on Political Stance Disaggregated by Gender
At the story-level, we first compare the no suggestions conditions to their corresponding liberal and conservative suggestions conditions. We see that adding conservative suggestions decreases the number of liberal characters. This trend is significant for “Thomas” ($t(123)=2.682$, $p_{\text{FDR}}\approx 0.0332$) and marginally significant for “Rebecca” ($t(112)=2.337$, $p_{\text{FDR}}\approx 0.0707$). We also see a marginal trend of “Thomas” being made conservative more often with conservative suggestions than without suggestions ($t(123)=-2.334$, $p_{\text{FDR}}\approx 0.0707$), with no such trend in the same setting for “Rebecca” ($t(112)=-0.984$, $p_{\text{FDR}}\approx 0.5407$). We see here that suggestions tend to successfully encourage participants to make characters liberal or conservative, but they are less successful in making “Rebecca” conservative, perhaps suggesting that participants have a harder time accepting suggestions of a fem-coded character being conservative.
We also compare rates of making characters liberal vs conservative within suggestion types. Without suggestions, we see no significant difference between making characters of any gender liberal or conservative (Rebecca: $t(66)=1.4$, $p_{\text{FDR}}\approx 0.3334$; Thomas: $t(78)=1.686$, $p_{\text{FDR}}\approx 0.2230$). We see “Thomas” is made liberal or conservative significantly more often depending on the direction of suggestions (conservative: $t(168)=-3.728$, $p_{\text{FDR}}\approx 0.0018$; liberal: $t(166)=3.315$, $p_{\text{FDR}}\approx 0.0066$). For “Rebecca”, we see significantly more liberal stories when they are suggested ($t(168)=4.575$, $p_{\text{FDR}}\approx 0.0001$), but the increase in conservative stories when they are suggested is not significant ($t(158)=-1.988$, $p_{\text{FDR}}\approx 0.1364$). This again shows that participants may have a harder time accepting suggestions of a fem-coded character being conservative, which is in agreement with (Cao et al., 2022b)’s findings about human perceptions of the political stance of women and supporting H1.
@@
We generally don’t see significant trends at the word level. We do see that participants accepted suggestions significantly more often in the condition where “Rebecca” is suggested to be liberal over “Thomas” (H2b; $t(3570)=-2.725$, $p_{\text{FDR}}\approx 0.0276$). This begins to suggest that participants are more comfortable with fem-coded characters being written as liberal than masc-coded ones, but this trend is not significant when we consider only words that specify the character’s stance.
Scenario Word Reliance Gender-Defining Word Reliance Gender-Defining Word Rejection Rate Story-Level Gender Freq.
@@
Teacher likeable When I was in school, Mr. Brown was my science teachers. He was very much witty and always try to make their lessons attractive with amazing experiments in such unique ways.
@@ A.5. Scenario Statistical Tests
Here, we include all statistical test and p-value tables for scenario-level experiments not otherwise included in the appendix. This includes overall story gender and other attribute rates (Table 6, Table 7, Table 8, and Table 9), overall word-level reliance (Table 11), reliance rates for attribute-specifying words (Table 12), and rejection rates for attribute-specifying suggestions (Table 13).
@@
M NS vs straight M $t(104)=-0.154$ $0.9479$
M NS vs straight F $t(104)=-0.556$ $0.7514$
M NS vs queer M $t(115)=-1.688$ $0.2218$
@@
M NS M vs F $t(56)=-8.825$ $0.0000$
M straight M vs F $t(152)=-16.063$ $0.0000$
M queer M vs F $t(174)=-6.035$ $0.0000$
F NS vs straight M $t(124)=-0.422$ $0.8234$
@@
F NS vs queer M $t(132)=2.973$ $0.0172$
F NS vs queer F $t(132)=-2.695$ $0.0328$
F NS M vs F $t(86)=7.704$ $0.0000$
F straight M vs F $t(162)=12.859$ $0.0000$
@@
NS vs M competitive competitive $t(121)=-1.297$ $0.3747$
NS vs M competitive unassertive $t(121)=-0.851$ $0.5921$
@@
NS vs M unassertive unassertive $t(117)=-3.024$ $0.0157$
NS vs F competitive competitive $t(125)=-3.464$ $0.0044$
@@
M competitive competitive vs unassertive $t(174)=3.898$ $0.0010$
M unassertive competitive vs unassertive $t(166)=-2.872$ $0.0207$
@@
F competitive competitive vs unassertive $t(174)=8.131$ $0.0000$
Student F unassertive competitive vs unassertive $t(156)=-0.883$ $0.5725$
NS vs M likeable repellent $t(129)=-0.123$ $0.9523$
NS vs M likeable likeable $t(129)=-2.489$ $0.0516$
NS vs M repellent repellent $t(120)=-1.48$ $0.2950$
NS vs M repellent likeable $t(120)=-0.778$ $0.6290$
NS vs F likeable repellent $t(115)=1.451$ $0.3065$
NS vs F likeable likeable $t(115)=-1.403$ $0.3291$
NS vs F repellent repellent $t(109)=-0.841$ $0.5980$
NS vs F repellent likeable $t(109)=1.261$ $0.3930$
M NS repellent vs likeable $t(74)=-3.195$ $0.0110$
M likeable repellent vs likeable $t(184)=-8.86$ $0.0000$
M repellent repellent vs likeable $t(166)=-3.774$ $0.0016$
F NS repellent vs likeable $t(68)=-4.747$ $0.0001$
F likeable repellent vs likeable $t(162)=-12.726$ $0.0000$
Teacher F repellent repellent vs likeable $t(150)=-3.886$ $0.0011$
NS vs M conservative conservative $t(123)=-2.334$ $0.0707$
NS vs M conservative liberal $t(123)=2.682$ $0.0332$
@@
NS vs M liberal liberal $t(122)=-0.39$ $0.8283$
NS vs F conservative conservative $t(112)=-0.984$ $0.5407$
@@
NS vs F liberal liberal $t(117)=-0.539$ $0.7605$
M NS conservative vs liberal $t(78)=-1.686$ $0.2230$
M conservative conservative vs liberal $t(168)=3.728$ $0.0018$
M liberal conservative vs liberal $t(166)=-3.315$ $0.0066$
F NS conservative vs liberal $t(66)=-1.4$ $0.3334$
@@
Town Hall F liberal conservative vs liberal $t(168)=-4.575$ $0.0001$
@@
untrustworthy NS vs F M $t(113)=4.238$ $0.0004$
untrustworthy NS vs M F $t(113)=-0.084$ $0.9634$
untrustworthy NS vs F F $t(113)=-4.367$ $0.0003$
trustworhty vs untrustworthy NS M $t(71)=-4.117$ $0.0008$
trustworhty vs untrustworthy M M $t(169)=-2.881$ $0.0206$
trustworhty vs untrustworthy F M $t(161)=-1.886$ $0.1667$
trustworhty vs untrustworthy NS F $t(71)=3.051$ $0.0161$
@@
F unconf vs F conf $t(4121)=-2.645$ $0.0331$
F unconf vs M unconf $t(4040)=-0.195$ $0.9213$
@@
Doctor F conf vs M conf $t(3650)=3.488$ $0.0031$
F vs M $t(6734)=-0.979$ $0.5407$
@@
M threatening vs M benevolent $t(3534)=-1.784$ $0.1927$
President F benevolent vs M benevolent $t(3545)=-1.56$ $0.2631$
@@
Wedding F straight vs M straight $t(3699)=-4.373$ $0.0001$
@@
F unassertive vs M unassertive $t(3615)=4.453$ $0.0001$
M unassertive vs M competitive $t(4135)=-6.49$ $0.0000$
Student F competitive vs M competitive $t(4188)=-2.0$ $0.1325$
F vs M $t(7091)=4.724$ $0.0000$
F untrustworthy vs F trustworthy $t(3443)=-3.168$ $0.0085$
@@
Detectives F trustworthy vs M trustworthy $t(3907)=5.235$ $0.0000$
F repellent vs F likeable $t(3591)=2.618$ $0.0351$

### v2

@@ A.2. Scenario: Student
In this scenario, the story prefix includes a team member with a traditionally feminine (“Abby”) or masculine (“John”) name. The model then suggests that this character is “competitive” or “unassertive” (See examples in Table 5). According to cao-etal-2022-theory, American annotators view men as comparatively more competitive and women as comparatively more unassertive though our participants may not have held this same stereotype (See Table 16).
@@ A.2.1. Effects on Competitiveness Disaggregated by Gender
When we consider competitiveness rates in overall stories, we see that suggestions often increase the respective rates. We see significantly more unassertive masc-coded (t(117)=3.024t(117)=-3.024, $p_{\text{FDR}}\approx 0.0157$, d=0.608d=-0.608), unassertive fem-coded (t(116)=2.874t(116)=-2.874, $p_{\text{FDR}}\approx 0.0210$, d=0.562d=-0.562), and competitive fem-coded students (t(125)=3.464t(125)=-3.464, $p_{\text{FDR}}\approx 0.0044$, d=0.666d=-0.666) when suggested when comparing to no suggestions conditions. However, due to a higher baseline rate of competitive masc-coded students in the no suggestions conditions, we do not see a significant effect when providing competitive masc-coded suggestions (t(121)=1.297t(121)=-1.297, $p_{\text{FDR}}\approx 0.3747$, d=0.259d=-0.259). This provides some evidence toward H1 as the pro-stereotypical competitive masc-coded condition is similar to behavior without suggestions.
We also see that generally providing suggestions increases competitiveness vs unassertiveness rates within a gender. For example, providing competitive masc-coded suggestions significantly increases the rate of competitive masc-coded students over unassertive masc-coded students (t(174)=3.898t(174)=-3.898, $p_{\text{FDR}}\approx 0.0010$, d=0.588d=-0.588). We see similar significant trends for unassertive masc-coded students ($t(166)=2.872$, $p_{\text{FDR}}\approx 0.0207$, $d=0.443$) and competitive fem-coded students (t(174)=8.131t(174)=-8.131, $p_{\text{FDR}}<0.0001$, d=1.226d=-1.226). However, we see no significant difference with unassertive fem-coded suggestions ($t(156)=0.883$, $p_{\text{FDR}}\approx 0.5725$, $d=0.141$). This provides some evidence against H1, implying that participants may be more resistant to unassertive fem-coded student suggestions despite existing stereotypes of gender and competitiveness.
@@
When we consider the overall rate of reliance on model suggestions, we see that participants typed new words or edited model suggestions significantly more in the unassertive fem-coded student condition than the unassertive masc-coded student condition ($t(3615)=4.453$, $p_{\text{FDR}}\approx 0.0001$, $d=0.148$) and similarity for the competitive masc-coded student condition over the unassertive masc-coded student condition (t(4135)=6.49t(4135)=-6.49, $p_{\text{FDR}}<0.0001$, d=0.202d=-0.202). These results do not support H2, but we see that none of these trends are significantly present when we constrain to words that determine the character’s competitiveness (H2a). When we consider model suggestions that affect competitiveness, we see marginally more acceptance of competitive fem-coded suggestions than unassertive (H2b; t(426)=2.159t(426)=-2.159, $p_{\text{FDR}}\approx 0.0999$, d=0.21d=-0.21).
Overall, the results in this scenario were mixed. We find a relatively low overall rate of clearly specifying the given student as competitive or unassertive which may have skewed the results. This may be due to poor scenario design where participants decided to focus on topics other than leadership or competitiveness, or it could be that in the classroom settings, participants held weaker internal stereotypes about gender and competitiveness than cao-etal-2022-theory’s held about more general settings. Indeed (as we show in Table 16) our post-survey results indicate that our participant pool may include more people who do not hold this stereotypical association between masculinity and competence (or, by extension, competitiveness).
@@ A.3. Scenario: Teachers
In this scenario, participants write about a teacher with a given gender where the model attempts to suggest the teacher’s personality as “likable” or “repellent” (See examples in Table 5). According to cao-etal-2022-theory, American annotators view men as comparatively less likable than women, and our participants indicated a significantly stronger association with straight women and warmth (which is associated with likability) than straight men (See Table 16).
A.3.1. Effects on Likability Disaggregated by Gender
Here, we discuss how suggestions affect the likability stance of “Mrs. Brown” and “Mr. Brown” (Table 4).
@@ A.3.1. Effects on Likability Disaggregated by Gender
Considering overall stories, we see that regardless of the teacher’s presumed gender, participants made the teacher likable significantly more often than repellent when not given suggestions (fem-coded teacher: t(68)=4.747t(68)=-4.747, $p_{\text{FDR}}\approx 0.0001$, d=1.135d=-1.135; masc-coded teacher: t(74)=3.195t(74)=-3.195, $p_{\text{FDR}}\approx 0.0110$, d=0.733d=-0.733). These trends continue to be significant regardless of suggestion type, including repellent suggestions (likable fem-coded: t(162)=12.726t(162)=-12.726, $p_{\text{FDR}}<0.0001$, d=1.987d=-1.987; repellent fem-coded: t(150)=3.886t(150)=-3.886, $p_{\text{FDR}}\approx 0.0011$, d=0.63d=-0.63; likable masc-coded: t(184)=8.86t(184)=-8.86, $p_{\text{FDR}}<0.0001$, d=1.299d=-1.299; repellent masc-coded: t(166)=3.774t(166)=-3.774, $p_{\text{FDR}}\approx 0.0016$, d=0.582d=-0.582). In other words, participants preferred to make the teacher likable, regardless of the presence or type of suggestions and regardless of the teacher’s gender as cued in the story prefix.
We also find that “Mr. Brown” is written as likable marginally more often with likable suggestions than without suggestions (t(129)=2.489t(129)=-2.489, $p_{\text{FDR}}\approx 0.0516$, d=0.479d=-0.479). However, perhaps due to a higher base rate of “Mrs. Brown” being written as likable, we see no such increase in likability with added likable suggestions for “Mrs. Brown” (t(115)=1.403t(115)=-1.403, $p_{\text{FDR}}\approx 0.3291$, d=0.283d=-0.283). In other words participants may have a stronger default preference for “Mrs. Brown” being likable, leading to a more limited effect of likable suggestions. This is in line with cao-etal-2022-theory’s and our findings about gender-likability stereotypes in humans. And these findings provide some support for H1 in that the proportion of likable “Mrs. Brown”s (pro-stereotypical) with no suggestions is more similar to the proportion of likable “Mrs. Brown”s with likable suggestions than the the proportion of likable “Mr. Brown”s (anti-stereotypical) with no suggestions is to the proportion of likable “Mr. Brown”s with likable suggestions.
@@
When we consider the rate of acceptances of model suggestions, we see significantly less acceptance of model suggestions in the condition where “Mrs. Brown” is suggested to be repellent over likable ($t(3591)=2.618$, $p_{\text{FDR}}\approx 0.0351$, $d=0.088$), while we see a significant effect in the opposite direction for “Mr. Brown” (t(3892)=3.162t(3892)=-3.162, $p_{\text{FDR}}\approx 0.0086$, d=0.102d=-0.102). We also see significantly less acceptance in the “Mr. Brown” is likable condition than for “Mrs. Brown” (t(4013)=5.922t(4013)=-5.922, $p_{\text{FDR}}<0.0001$, d=0.187d=-0.187). These results support H2, as we can see that more pro-stereotypical conditions (e.g., suggesting a fem-coded character is likable) lead to more acceptance of suggestions.
However, the effects are quite different when we only consider words that determine likability. Here we see a trend of participants accepting more “likable” suggestions over “repellent” for either gender, though the effect is only significant for “Mr. Brown” (masc-coded: $t(179)=2.673$, $p_{\text{FDR}}\approx 0.0331$, $d=0.398$; fem-coded: $t(157)=2.068$, $p_{\text{FDR}}\approx 0.1190$, $d=0.328$). Under H2a, we would have expected to see less acceptance of “likable” suggestions for the masc-coded “Mr. Brown”. When we consider the acceptance rate of model suggestions, we see higher rates of acceptance of “likable” suggestions for either gender, but in this case, it is only significant for “Mrs. Brown” (masc-coded: t(438)=2.446t(438)=-2.446, $p_{\text{FDR}}\approx 0.0533$, d=0.233d=-0.233, $p_{\text{FDR}}\approx 0.0331$; fem-coded: t(383)=4.106t(383)=-4.106, $p_{\text{FDR}}\approx 0.0004$, d=0.42d=-0.42), supporting H2b.
These results show that participants may have preferred suggestions of teachers of any gender being likable. This seems reasonable as at the story-level likable teachers were generally preferred even without suggestions. While these results do not match our overall hypotheses about reliance under pairs of gender and likability, they may suggest that participants’ stereotypes about teachers being likable people were stronger than their stereotypes about people of different genders being likable.
@@ A.4. Scenario: Town Hall
In this scenario, the story prefix includes a town hall participant with a traditionally feminine (“Rebecca”) or masculine (“Thomas”) name. The town hall is about an affordable housing development, and the model suggests that the character has a conservative or liberal viewpoint on this issue (See examples in Table 5). According to cao-etal-2022-theory, American annotators view men as comparatively more conservative than women, an association echoed by our participants in our post-survey (See Table 16).
@@ A.4.1. Effects on Political Stance Disaggregated by Gender
At the story-level, we first compare the no suggestions conditions to their corresponding liberal and conservative suggestions conditions. We see that adding conservative suggestions decreases the number of liberal characters. This trend is significant for “Thomas” ($t(123)=2.682$, $p_{\text{FDR}}\approx 0.0332$, $d=0.514$) and marginally significant for “Rebecca” ($t(112)=2.337$, $p_{\text{FDR}}\approx 0.0707$, $d=0.478$). We also see a marginal trend of “Thomas” being made conservative more often with conservative suggestions than without suggestions (t(123)=2.334t(123)=-2.334, $p_{\text{FDR}}\approx 0.0707$, d=0.447d=-0.447), with no such trend in the same setting for “Rebecca” (t(112)=0.984t(112)=-0.984, $p_{\text{FDR}}\approx 0.5407$, d=0.201d=-0.201). We see here that suggestions tend to successfully encourage participants to make characters liberal or conservative, but they are less successful in making “Rebecca” conservative, perhaps suggesting that participants have a harder time accepting suggestions of a fem-coded character being conservative.
We also compare rates of making characters liberal vs conservative within suggestion types. Without suggestions, we see no significant difference between making characters of any gender liberal or conservative (Rebecca: $t(66)=1.4$, $p_{\text{FDR}}\approx 0.3334$, $d=0.340$; Thomas: $t(78)=1.686$, $p_{\text{FDR}}\approx 0.2230$, $d=0.377$). We see “Thomas” is made liberal or conservative significantly more often depending on the direction of suggestions (conservative: t(168)=3.728t(168)=-3.728, $p_{\text{FDR}}\approx 0.0018$, d=0.571d=-0.571; liberal: $t(166)=3.315$, $p_{\text{FDR}}\approx 0.0066$, $d=0.511$). For “Rebecca”, we see significantly more liberal stories when they are suggested ($t(168)=4.575$, $p_{\text{FDR}}<0.0001$, $d=0.701$), but the increase in conservative stories when they are suggested is not significant (t(158)=1.988t(158)=-1.988, $p_{\text{FDR}}\approx 0.1364$, d=0.314d=-0.314). This again shows that participants may have a harder time accepting suggestions of a fem-coded character being conservative, which is in agreement with (cao-etal-2022-theory)’s and our findings about human perceptions of the political stance of women and supporting H1.
@@
We generally don’t see significant trends at the word level. We do see that participants accepted suggestions significantly more often in the condition where “Rebecca” is suggested to be liberal over “Thomas” (H2b; t(3570)=2.725t(3570)=-2.725, $p_{\text{FDR}}\approx 0.0276$, d=0.091d=-0.091). This begins to suggest that participants are more comfortable with fem-coded characters being written as liberal than masc-coded ones, but this trend is not significant when we consider only words that specify the character’s stance.
Scenario Word Reliance
Gender-Defining Word
Reliance
Gender-Defining Word
Rejection Rate
Story-Level
Gender Freq.
@@
Teacher likable When I was in school, Mr. Brown was my science teachers. He was very much witty and always try to make their lessons attractive with amazing experiments in such unique ways.
@@ A.5. Scenario Statistical Tests
Here, we include all statistical test and p-value tables for scenario-level experiments not otherwise included in the appendix. This includes overall story gender and other attribute rates (Table 6, Table 7, Table 8, and Table 9), overall word-level reliance (Table 11), reliance rates for attribute-specifying words (Table 12), and rejection rates for attribute-specifying suggestions (Table 13). For details of the process obtaining the relevant outcome measurements using LLM annotation, please see subsection 5.2.
@@
M NS vs straight M t(104)=0.154t(104)=-0.154 $0.9479$
M NS vs straight F t(104)=0.556t(104)=-0.556 $0.7514$
M NS vs queer M t(115)=1.688t(115)=-1.688 $0.2218$
@@
M NS M vs F t(56)=8.825t(56)=-8.825 $0.0000$ *
M straight M vs F t(152)=16.063t(152)=-16.063 $0.0000$ *
M queer M vs F t(174)=6.035t(174)=-6.035 $0.0000$ *
F NS vs straight M t(124)=0.422t(124)=-0.422 $0.8234$
@@
F NS vs queer M $t(132)=2.973$ $0.0172$ *
F NS vs queer F t(132)=2.695t(132)=-2.695 $0.0328$ *
F NS M vs F $t(86)=7.704$ $0.0000$ *
F straight M vs F $t(162)=12.859$ $0.0000$ *
@@
NS vs M competitive competitive t(121)=1.297t(121)=-1.297 $0.3747$
NS vs M competitive unassertive t(121)=0.851t(121)=-0.851 $0.5921$
@@
NS vs M unassertive unassertive t(117)=3.024t(117)=-3.024 $0.0157$ *
NS vs F competitive competitive t(125)=3.464t(125)=-3.464 $0.0044$ *
@@
M competitive competitive vs unassertive $t(174)=3.898$ $0.0010$ *
M unassertive competitive vs unassertive t(166)=2.872t(166)=-2.872 $0.0207$ *
@@
F competitive competitive vs unassertive $t(174)=8.131$ $0.0000$ *
Student F unassertive competitive vs unassertive t(156)=0.883t(156)=-0.883 $0.5725$
NS vs M likable repellent t(129)=0.123t(129)=-0.123 $0.9523$
NS vs M likable likable t(129)=2.489t(129)=-2.489 $0.0516$
NS vs M repellent repellent t(120)=1.48t(120)=-1.48 $0.2950$
NS vs M repellent likable t(120)=0.778t(120)=-0.778 $0.6290$
NS vs F likable repellent $t(115)=1.451$ $0.3065$
NS vs F likable likable t(115)=1.403t(115)=-1.403 $0.3291$
NS vs F repellent repellent t(109)=0.841t(109)=-0.841 $0.5980$
NS vs F repellent likable $t(109)=1.261$ $0.3930$
M NS repellent vs likable t(74)=3.195t(74)=-3.195 $0.0110$ *
M likable repellent vs likable t(184)=8.86t(184)=-8.86 $0.0000$ *
M repellent repellent vs likable t(166)=3.774t(166)=-3.774 $0.0016$ *
F NS repellent vs likable t(68)=4.747t(68)=-4.747 $0.0001$ *
F likable repellent vs likable t(162)=12.726t(162)=-12.726 $0.0000$ *
Teacher F repellent repellent vs likable t(150)=3.886t(150)=-3.886 $0.0011$ *
NS vs M conservative conservative t(123)=2.334t(123)=-2.334 $0.0707$
NS vs M conservative liberal $t(123)=2.682$ $0.0332$ *
@@
NS vs M liberal liberal t(122)=0.39t(122)=-0.39 $0.8283$
NS vs F conservative conservative t(112)=0.984t(112)=-0.984 $0.5407$
@@
NS vs F liberal liberal t(117)=0.539t(117)=-0.539 $0.7605$
M NS conservative vs liberal t(78)=1.686t(78)=-1.686 $0.2230$
M conservative conservative vs liberal $t(168)=3.728$ $0.0018$ *
M liberal conservative vs liberal t(166)=3.315t(166)=-3.315 $0.0066$ *
F NS conservative vs liberal t(66)=1.4t(66)=-1.4 $0.3334$
@@
Town Hall F liberal conservative vs liberal t(168)=4.575t(168)=-4.575 $0.0001$ *
@@
untrustworthy NS vs F M $t(113)=4.238$ $0.0004$ *
untrustworthy NS vs M F t(113)=0.084t(113)=-0.084 $0.9634$
untrustworthy NS vs F F t(113)=4.367t(113)=-4.367 $0.0003$ *
trustworhty vs untrustworthy NS M t(71)=4.117t(71)=-4.117 $0.0008$ *
trustworhty vs untrustworthy M M t(169)=2.881t(169)=-2.881 $0.0206$ *
trustworhty vs untrustworthy F M t(161)=1.886t(161)=-1.886 $0.1667$
trustworhty vs untrustworthy NS F $t(71)=3.051$ $0.0161$ *
@@
F unconf vs F conf t(4121)=2.645t(4121)=-2.645 $0.0331$ *
F unconf vs M unconf t(4040)=0.195t(4040)=-0.195 $0.9213$
@@
Doctor F conf vs M conf $t(3650)=3.488$ $0.0031$ *
F vs M t(6734)=0.979t(6734)=-0.979 $0.5407$
@@
M threatening vs M benevolent t(3534)=1.784t(3534)=-1.784 $0.1927$
President F benevolent vs M benevolent t(3545)=1.56t(3545)=-1.56 $0.2631$
@@
Wedding F straight vs M straight t(3699)=4.373t(3699)=-4.373 $0.0001$ *
@@
F unassertive vs M unassertive $t(3615)=4.453$ $0.0001$ *
M unassertive vs M competitive t(4135)=6.49t(4135)=-6.49 $0.0000$ *
Student F competitive vs M competitive t(4188)=2.0t(4188)=-2.0 $0.1325$
F vs M $t(7091)=4.724$ $0.0000$ *
F untrustworthy vs F trustworthy t(3443)=3.168t(3443)=-3.168 $0.0085$ *
@@
Detectives F trustworthy vs M trustworthy $t(3907)=5.235$ $0.0000$ *
F repellent vs F likable $t(3591)=2.618$ $0.0351$ *